What is the value of the equilibrium constant for
Answer #1

G = -RT*lnK

since K = Ksp

then

G = -8.314*298*ln(K)

Grxn = -nF*E°cell

Al3+(aq) + 3e– → Al(s) –1.66

Cd2+(aq) + 2e– → Cd(s) –0.403

E°cell = Ered - Eox = -0.403 + 1.66 = 1.257V

GRxn = -6*96500*(1.257) = -727803 J/molK

G = -8.314*298*ln(K)

-727803 =  -8.314*298*ln(K)

K = exp(727803/(8.314*298)) = 3.7744*10^127